OpenStudy (anonymous):

can any one help me in this Exercises Find any horizontal and vertical asymptotes and any holes that may exist for each rational function y = x^2+7x+12 / x+4

OpenStudy (anonymous):

y = x^2+7x+12 y= (x+3)(x+4) \[y=\frac{(x+3)(x+4)}{x+4}\]x+4 cancel\[y= x+3\]

OpenStudy (anonymous):

no asymptotes or holes , it slope intercept

OpenStudy (anonymous):

y intercept = 3